OpenStudy (anonymous):

How many solutions are there to the equation y = 14x – 9? all i want is an example of how to answer this problem! Will medal and fan

OpenStudy (anonymous):

Please help!

Directrix (directrix):

y = 14x – 9 is the equation of a line. A line has infinitely many points on it. On the coordinate plane, each of these points has an ordered pair (x,y) to name it.

Directrix (directrix):

@ChiShiki You could argue infinitely many in this way: y = 14x – 9 --------- Let x = 1, then y = 14*1 -9 = 5 which is one solution Let x = 2, then y = 14*2 - 9 = 19 which is another solution Let x = whatever and then crank out the y for that x

Directrix (directrix):

Question?

OpenStudy (anonymous):

Thank You so much!